OpenStudy (anonymous):

In 1977, the price of a scientific calculator was $175. In 2004, the price was $15. Assuming the change in price was linear, what was the approximate price of a scientific calculator in 1998? * help please*

OpenStudy (anonymous):

In 1977, price= 175 dollars In 2004, price=15 dollars Difference in no. of years=2004-1977=27 years Change in price= 175-15=160 Or, in 27 years, price reduces by 160 dollars. Now, 1998-1977=21 years Therefore, in 21 years, reduction in price=(160/27)*21=124.45 Dollars (approx)

OpenStudy (anonymous):

thanks darlling (: